if x - 4, 4 - x and x form an arithmetic progression, what is x? please help with a little explanation.thanks in advance.
Answer by Boreal(15235)

(x-4),(4-x),x must have equal differences between them
Difference between first and second is (4-x)-(x-4)=4-2x+4=8-2x
The difference between second and third must equal 8-2x. It is x-(4-x)=2x-4
Therefore, 8-2x=2x-4
12=4x by adding 2x and 4 to both sides;
x=3 ANSWER
-1, 1, 3 are the numbers, and they form an arithmetic progression with difference 2 between each.